QUANTUM PHYSICS You have accidentally been beamed into an alternative universe, where first graders love broccoli, where highTwo skateboarders push off from each other as we did in lecture. One has a mass of 80 kg, and the other has a mass of 72 kg.Two skateboarders push off from each other as we did in lecture. One has a mass of 80 kg, and the other has a mass of 72 kg.We say that the momenta exchanged by two skateboarders are equal in size because we know that (i) the interaction time At is
Answer #1

Wavelength of the neutron P, = 250000m 3.6 1.44 x 10-5 in = 250km F F me V V2 When two skateboarders of different masses pushTherefore the skateboarders exchange equal but opposite amounts of momentum Now -4 P = AP2 -m4 vi = m 24 V2 Since m 7 m2 A vi